Orion Corporation has established the following standards for the prime costs of one unit of its chief product, dartboards.

Standard Quantity Standard Price or Rate Standard Cost

Direct material 6.00pounds $1.10per pound  $6.60  

Direct labor 0.30hour $6.00per hour   1.80  

Total         $8.40  

During June, Orion purchased 227,000 pounds of direct material at a total cost of $274,670. The total wages for June were $72,576, 75 percent of which were for direct labor. Orion manufactured 36,000 dartboards during June, using 205,200 pounds of the direct material purchased in June and 11,520 direct-labor hours.

Required:

Compute the following variances for June. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "Favorable" or "Unfavorable". Select "None" and enter "0" for no effect (i.e., zero variance).

Calculate 1. Direct Material Price Variance, 2. Direct Material quantity Variance, 3. Direct material purchase price variance, 4. Direct labor rate variance, 5. Direct-labor efficiency variance
Answer #1

Details for Standard & Actual Quantity of Material & Hours for 36000 units

Particulars Standard Quantity Standard Rate Actual Quantity Actual Rate
Direct Material 6*36000=216000 units $1.10 205,000 Units 274,670/227,000 = $1.30
Direct Labor 0.30*36000=10,800 hours $6.00 11,520 $72,576/(11520/75%) = 4.725

On the basis of the above data, we had calculated the following variance

1. Direct Material Price Variance =  Actual Quantity * (Standard Price - Actual Price)

= 205,000 * (1.10-1.30) = -41,000 (Unfavorable)

2. Direct Material Price Variance = Standard Price * (Standard Quantity - Actual Quantity)

= 1.1 * (216,000 - 205,000) = 12,100 (favorable)

3. Direct Material Purchase Price Variance = Actual Quantity * (Actual Unit Cost - Standard Unit cost)

= 205,000 * (1.3 - 1.1) = 41,000

4. Direct Labor Rate Variance = Actual Hour * (Standard Rate - Actual Rate)

= 11,520 * (6.000 - 4.725) = 14,688 (Favorable)

5. Direct Labor Efficiency Variance = Standard Rate * (Actual Hours - Standard Hours)

= 6.00 * (11,520 - 10,800) = 4320 (Favorable)